SLAC PEP-II
BABAR
SLAC<->RAL
Babar logo
HEPIC E,S & H Databases PDG HEP preprints
Organization Detector Computing Physics Documentation
Personnel Glossary Sitemap Search Hypernews
Unwrap page!
Det. Search
Who's who?
Meetings
FAQ
Images
Archive
Systems
Performance
Intern. region
Vertex Tracker
Drift chamber
DIRC
Calorimeter
IFR
LST
Magnet
Electronics
Trigger
Operations
Run Coordination
Contact Experts
Shift Takers Info
Operations Manual
Electronic Logbook
Ops Hypernews
Shift Signup
Check this page for HTML 4.01 Transitional compliance with the
W3C Validator
(More checks...)

Trigger Operation Meeting (11/13/07)

Attendees: Rainer Bartoldus, Su Dong, Mark Tibbetts, Rahmat Rahmat, Al Eisner, Debbie Bard, Haiwen Zhao, Kim Hojeong.

Meeting Information

Date : November 13, 2007 Time : 2:00pm - 3:30pm PST SLAC Room : Madrone conf room, ROB Dial-in : 1-510-665-5437, pass code 6649#

--- Minutes ---

Operations

  • All crates are kept powerd on due to cosmics runs from time to time.

EMT

  • Mark: EMT code change test. not successful. Problem with script language.

DCT

GLT

Online

  • Fang: tried to do DCT-DCZ calibration at test stand but doesn't work. She'll try it in IR2 to see if it works there, and try again at test stand. She'll also update the instruction if needed.

L3

  • Su Dong: has some plots on L3EMT filter study. (current plots, previous plots and discussion: look for L3 first bullet.).
    • When cut on L3 mass cut at 800 MeV, not good enough to see f0(980). (plot)
    • Try at 650 MeV, looks better. (plot)
    • Al thinks the distribution (mass of pi0pi0 vs. Reco EMC mass) looks weird. Maybe it is flat in m^2 but we are looking it in m.
    • Su Dong will run some test for rate.
    • Maybe apply missing pT cut? (Al) Su Dong thinks the best combination could be pT cut for low mass and no pT cut for high mass.
    • Rainer: maybe relaxing an exisiting filter cut (1.5GeV) and another filter may work too.

DQM

  • Rahmat: data quality is fine. Working on strip chart for L4 quantities. Trouble getting L3 lumi from ir2runs. The numbers should be in L3 histograms (Rainer).
  • Hojeong: JAS page updated. EMT page will be updated with link from Mark on how to find bins.

Trickle injection monitoring

  • Al: Single Run, Multiple Runs.
    • He's ready to check in, comments welcome.
    • Statistics is low for each run, but will put for each run for now. Later, maybe accumulate 10 runs or so for high statistics.
    • The first page looks weird on screen, hard to see anything. But it looks fine if printed.
    • The shadow area is hard coded! Showing the numbers in the second page.
    • There is an edge effect. It shows low stat in the first bin.
    • The third page has line at 1. Maybe he'll put fit either poly0 or poly1.
    • The multiple run shows structure due to deadtime.

Trigger Configuration

  • Hojeong: another emulation with new order is ready for all possible modes. (Survey page) She will make emulation2 on L1passthru run.
  • Al: 3B&2A, M*&1Z lines are used for physics. Rainer & Su Dong: they are not. Al will show summary of his study by next meeting.
  • Haiwen: made a summary page
  • Need to ask analysts to DO put their analysis cuts (including fiducial cuts to make sure we are looking at the events people are interested) and DO NOT put any trigger or filter requirements. Many skims require trigger and filter, so NO SKIMS.
  • FYI:
    • Run1-4: no Z, no Zk
    • Run5: yes Z, no Zk
    • Run6: yes Z, yes Zk

AOB

  • We will have meeting next week.

This page was created by Kim Hojeong; however, you are welcome to edit and/or correct above items, specially for those who attended the meeting. Thanks.

Last Modified: November 13, 2007.